OpenStudy (anonymous):

A line has a slope of 9/4 and a y-intercept of -14. Which equations represent the line? Choose ALL answers that are correct. A. -9x + 4y = -56 B. 9x + 4y = 56 C. y = 9/4x - 14 D. y = 9/4x + 14

OpenStudy (texaschic101):

y = 9/4x - 14 <=== this one y = 9/4x - 14 --- multiply entire equation by common denominator of 4 4y = 9x - 56 --- subtract 9x from both sides -9x + 4y = -56 <=== this one 9x - 4y = 56 <== this would also be one, but it is not an answer choice so, A and C

OpenStudy (anonymous):

Thanks! Could you help with another?